OpenStudy (anonymous):

If C = 52 degrees , a = 25, c = 22, solve the triangle Find A. (If there is more than one solution, find the larger value of A.)

OpenStudy (anonymous):

a/ sin A = c / sin C

OpenStudy (anonymous):

a/ sin A = c / sin C 25/sin A =22/sin 52 sin A = (25/22) sin 52

OpenStudy (anonymous):

sin A = (25/22) sin 52 sin A = (25/22) 0.788 sin A = 0.8954 A = sin^-1 (0.8954) A = 63.5 degrees

OpenStudy (anonymous):

I understand what you did, but once I put in my the answer it came out wrong. Is 63.5 rounded off of an other number?

OpenStudy (anonymous):

try 63.56...,

OpenStudy (anonymous):

No never mind, I put in the full number it didn't work here let me provide the triangle

OpenStudy (anonymous):